# A Watermark for Low-entropy and Unbiased Generation in Large Language Models

### v1

@@ 5.3 Results on HumanEval
In this section, we compare our methods against baselines on the HumanEval dataset. Table 2 presents the perplexity, pass@$k$ scores, and watermark strength for all methods. First, we focus on the result analysis for all unbiased watermarks. As reported, our STA-1 method achieves similar perplexity and pass@$k$ scores compared to no watermarking and other unbiased watermarking methods.
@@
Moreover, we examine the risk of unsatisfactory outputs produced by unbiased watermarks for low-entropy generations. Specifically, we compare different unbiased watermarks in terms of four more metrics. We ran 10 times of code generation for each problem using different unbiased watermarking methods with 10 different keys. Table 3 reports the variance of perplexity, the number of passed problems, [^2] If the problem is solved by any one generation out of 10 runs, it is considered passed. the number of passed codes, and the average number of passed codes among all passed problems. In particular, the STA-1 method demonstrates the lowest variance of perplexity compared to Dipmark($\alpha$=0.3), Dipmark($\alpha$=0.4), and $\gamma$-reweight with a variance of 1.461 compared to 1.535, 1.853, and 1.722, respectively. A lower variance indicates a lower risk among different text generations under different keys. Additionally, we show the average number of passed codes among all passed problems. For example, 3.136 in Table 3 means among all solved problems, an average of 3.136 generated codes are accurate w.r.t. 10 generations by STA-1. We conclude from Table 3 that although Dipmark solves more problems, it fails to provide consistent accurate codes among different generations. Instead, our method outperforms other unbiased watermarks (Dipmark($\alpha$=0.3), Dipmark($\alpha$=0.4), $\gamma$-reweight) in providing consistency, with an average number of passed codes of 3.136 compared to 2.675, 2.631, and 2.774, respectively. In summary, the STA-1 method has a lower risk when generating low-entropy texts, as discussed in Theorem 3.
In terms of watermark strength, STA-M ($M\in\{4,8,16\}$) yields higher watermark strength in comparison to all unbiased watermarks while maintaining similar pass scores. The STA-16 method achieves comparable watermark strength against biased watermark KGW($\delta=2$) with an AUC of 0.764 ($z=2$) against 0.795. The text quality is maintained with a pass@$10$ of 0.476, highlighting the efficacy of the heuristics to enhance watermark strength at high-entropy generation steps.
@@
        Attack Setting No Attack GPT-3.5 DIPPER-1 DIPPER-2 Copy-Paste
@@
KGW($\delta=1$) 0.96 0.96 0.27 0.57 0.13 0.53 0.15 0.54 0.68 0.75
KGW($\delta=1.5$) 0.98 0.98 0.41 0.62 0.22 0.56 0.27 0.57 0.90 0.90
KGW($\delta=2$) 0.98 0.98 0.54 0.68 0.30 0.58 0.40 0.62 0.95 0.95
Dipmark($\alpha=0.3$) 0.92 0.92 0.29 0.57 0.24 0.55 0.26 0.55 0.61 0.70
Dipmark($\alpha=0.4$) 0.95 0.95 0.38 0.61 0.31 0.58 0.34 0.59 0.75 0.79
$\gamma$-reweight 0.97 0.97 0.41 0.61 0.32 0.57 0.36 0.60 0.74 0.78
STA-1 0.96 0.96 0.47 0.63 0.39 0.60 0.46 0.63 0.78 0.81
STA-4($\tau$=1.35) 0.98 0.98 0.72 0.78 0.65 0.73 0.69 0.75 0.95 0.95
STA-8($\tau$=1.35) 0.97 0.97 0.78 0.81 0.71 0.77 0.76 0.79 0.95 0.95
STA-16($\tau$=1.35) 0.97 0.97 0.76 0.80 0.68 0.74 0.78 0.81 0.95 0.95
@@ 5.4 Attacking STA
We assessed the robustness of different watermarking methods under different attacks consisting of paraphrasing using GPT-3.5, two configurations of the DIPPER attack [14], and the copy-paste attack [12]. Detailed settings of different attacks are described in Appendix E.4. Table 4 reports the F1-score and AUC of watermark detection under each attack with $z=2$. As reported, STA-M is robust against different attacks. The reason is that GPT-3.5 and DIPPER are both LLM-based attacks that sample tokens to replace tokens in given texts. However, LLM-based attacks can hardly replace tokens in STA-M generated text because of the high text quality, and cannot find enough tokens to replace them because of the high watermark strength of STA-M. Meanwhile, the high watermark strength ensures robustness against the copy-paste attack.
@@ 6 Conclusions and Future Work
In this work, we propose a new unbiased watermarking method named STA-1. We clarify the text quality (regarding the risk of unsatisfactory outputs) and watermark strength tradeoff of unbiased watermarks in low-entropy scenarios. We also extend STA-1 to STA-M which can enhance watermark strength with little text quality shifts. Experimental results on low-entropy datasets prove that STA-1 is comparable to other unbiased watermarks and has a low risk. Moreover, the results from both datasets demonstrate the efficacy and robustness of STA-M. Future work of our study can be categorized into several ways. First, watermarking low-entropy tasks is still challenging and future work can devise better watermarking methods. Second, future work could incorporate more datasets and generative LLMs for evaluation of our method.
@@ 7 Related Work
With the development of LLMs, the idea of watermarking LLMs has been proposed [1, 12] and widely explored [18]. Existing white-box watermarking techniques can be categorized into watermarking during logits and probabilities generation [13, 16, 11, 30, 8, 35, 34, 25, 28], and watermarking by sampling candidate tokens [5, 15, 10]. We refer readers to Appendix F for a detailed discussion on related work.
Appendix A Details of Previous Methods
@@
KGW [12] first randomly splits the vocabulary set $\mathcal{V}$ into two non-overlapping lists based on a uniformly distributed key $k$: a ‘green’ list and a ‘red’ list. This method has two versions: the ‘hard’ version completely ignores the red list tokens and only samples tokens from the green list; The ‘soft’ version adds a predefined constant $\delta$ to logits of green list tokens while keeping logits of red list tokens fixed. The soft KGW reweights distribution as
@@
Wu et al. [33] propose an unbiased reweighting method, named Dipmark. Dipmark arranges all probability masses over the vocabulary set from the original LLM output consecutively within the interval $[0,1]$ and then randomly permutes their orders based on a key $k$. A hyperparameter $\alpha\in[0,0.5]$ partitions the probability interval $[0,1]$ into three segments: $[0,\alpha]$, $(\alpha,1-\alpha]$, and $(1-\alpha,1]$. Probability masses in the first segment are set to 0, those in the second remain constant, and those in the third are doubled. Denote the token order after permutation as $\widetilde{\mathcal{V}}$, the adjusted probability for the $j$-th token within $\widetilde{\mathcal{V}}$ is $P_{M,w}(x^{t}=j|x^{-N_{p}:(t-1)};k)=F(j|\widetilde{\mathcal{V}})-F(j-1|\widetilde{\mathcal{V}})$, with $F(j|\widetilde{\mathcal{V}})$ being defined as
@@
Appendix B Proofs
B.1 Proof of Theorem 1
@@
B.2 Proof of Theorem 2
In this proof, we employ the notations introduced in the proof of Theorem 1 in Section B.1, and we leverage the results derived from that theorem’s proof.
@@ B.2 Proof of Theorem 2
B.3 Proof of Theorem 3
In this proof, we continue utilizing the notations introduced in the proof of Theorem 1 in Section B.1.
@@ B.3 Proof of Theorem 3
$\mathbb{V}_{k\sim P_{K}(k)}^{\text{STA-1}}\left[p_{max}^{w,k}\right]<(1-p_{max})^{2}\frac{(N+1)}{3(N-1)}=\mathbb{V}_{k\sim P_{K}(k)}^{\text{Dipmark}}\left[p_{max}^{w,k}\right],$
@@
Appendix C Example of Risk-averse
St. Petersburg paradox [31]. Assume that one must choose either one lottery from the following two lotteries. (1) Lottery 1 ($L1$) has a 0.8 probability of earning nothing and the other 0.2 probability of losing 1,000 dollars. (2) Lottery 2 ($L2$) has a 0.5 probability of losing 100 dollars and the other 0.5 probability of losing 300 dollars.
@@ Appendix C Example of Risk-averse
Computationally, assume the person has 1,001 dollars in total and the utility function is $\ln(Y)$ [6], where $Y$ is the wealth. The utility function measures happiness. It is a concave function (such as $\ln(Y)$) because people are happier if they are wealthier ($\ln^{\prime}(Y)>0$) but the increment of happiness decreases as the wealth increases ($\ln^{\prime\prime}(Y)<0$).
@@ Appendix D STA-M Details
Appendix E Experiment
E.1 Experimental Setup
Datasets and metrics. We employed two public datasets which are C4 subset [24, 12] for news-like text generation and HumanEval [4] for code generation. Specifically, C4 represents the high-entropy generation task and HumanEval represents the low-entropy generation task.
C4: We extracted random text segments from the news-like subset of the C4 dataset [24] following Kirchenbauer et al. [12]. For each segment, we removed a fixed number of tokens from the end and the removed tokens served as a ‘baseline’ completion. The remaining tokens were used as the prompt.

### v2

@@ 5.3 Results on HumanEval
In this section, we compare our methods against baselines on the HumanEval dataset. Table 2 presents the perplexity, pass@$k$ scores, and watermark strength for all methods. Since it is better not to control the length of a code during generation, we remove detection time results. First, we focus on the result analysis for all unbiased watermarks. As reported, our STA-1 method achieves similar perplexity and pass@$k$ scores compared to no watermarking and other unbiased watermarking methods.
@@
Moreover, we examine the risk of unsatisfactory outputs produced by unbiased watermarks for low-entropy generations. Specifically, we compare different unbiased watermarks in terms of four more metrics. We ran 10 times of code generation for each problem using different unbiased watermarking methods with 10 different keys. Table 3 reports the average variance of perplexity among each problem, the number of passed problems (if the problem is solved by any one generation out of 10 runs, it is considered passed), the number of passed codes, and the average number of passed codes among all passed problems. In particular, the STA-1 method demonstrates the lowest variance of perplexity compared to RDW, Dipmark($\alpha$=0.3), Dipmark($\alpha$=0.4), and $\gamma$-reweight with a variance of 1.461 compared to 2.202, 1.535, 1.853, and 1.722, respectively. A lower variance indicates a lower risk among different text generations under different keys. Additionally, we show the average number of passed codes among all passed problems. For example, 3.136 in Table 3 means among all solved problems, an average of 3.136 generated codes are accurate w.r.t. 10 generations by STA-1. We conclude from Table 3 that although Dipmark solves more problems, it fails to provide consistent accurate codes among different generations. Instead, our method outperforms other unbiased watermarks (RDW, Dipmark($\alpha$=0.3), Dipmark($\alpha$=0.4), $\gamma$-reweight) in providing consistency, with an average number of passed codes of 3.136 compared to 2.844, 2.675, 2.631, and 2.774, respectively. In summary, the STA-1 method has a lower risk when generating low-entropy texts, as discussed in Theorem 3.
In terms of watermark strength in Table 2, STA-M ($M\in\{4,8,16\}$) yields higher watermark strength in comparison to all unbiased watermarks while maintaining similar pass scores. The STA-16 method achieves comparable watermark strength against biased watermark KGW($\delta=2$) with an AUC of 0.764 ($z=2$) against 0.795. The text quality is maintained with a pass@$10$ of 0.476, highlighting the efficacy of the heuristics to enhance watermark strength at high-entropy generation steps.
@@
        Attack Setting No Attack Copy-Paste GPT-3.5 DIPPER-1 DIPPER-2
@@
KGW($\delta=1$) 0.96 0.96 0.68 0.75 0.27 0.57 0.13 0.53 0.15 0.54
KGW($\delta=1.5$) 0.99 0 98 0.90 0.90 0.41 0.62 0.22 0.56 0.27 0.57
KGW($\delta=2$) 0.99 0.99 0.95 0.95 0.54 0.68 0.30 0.58 0.40 0.62
RDW 0.98 0.98 0.83 0.79 0.73 0.78 0.64 0.73 0.65 0.73
Dipmark($\alpha=0.3$) 0.93 0.94 0.61 0.70 0.29 0.57 0.24 0.55 0.26 0.55
Dipmark($\alpha=0.4$) 0.96 0.96 0.75 0.79 0.38 0.61 0.31 0.58 0.34 0.59
$\gamma$-reweight 0.96 0.96 0.74 0.78 0.41 0.61 0.32 0.57 0.36 0.60
STA-1 0.96 0.96 0.78 0.81 0.47 0.63 0.39 0.60 0.46 0.63
STA-4($\tau$=1.35) 0.97 0.97 0.95 0.95 0.72 0.78 0.65 0.73 0.69 0.75
STA-8($\tau$=1.35) 0.98 0.98 0.95 0.95 0.78 0.81 0.71 0.77 0.76 0.79
STA-16($\tau$=1.35) 0.97 0.97 0.95 0.95 0.76 0.80 0.68 0.74 0.78 0.81
@@ 5.4 Attacking STA
We assessed the robustness of different watermarking methods under different attacks consisting of the copy-paste attack (Kirchenbauer et al., 2023a), paraphrasing using GPT-3.5, and two configurations of the DIPPER attack (Krishna et al., 2024). Detailed settings of different attacks are described in Appendix F.4.
Table 4 reports the F1-score and AUC of watermark detection under each attack with $z=2$. As reported, for unbiased watermarks, RDW achieves the best result since its detection framework based on brute force search is designed to solve the robustness issue (Kuditipudi et al., 2023). In contrast, STA-M is robust against different attacks with a low detection time. For the copy-paste attack, since STA-M is based on the green-red list partition and changing a token can only affect the detection score of itself and the next token, it is naturally robust to simple text insertion and removal (Kirchenbauer et al., 2023a). Meanwhile, LLM-based attacks, such as GPT-3.5 and DIPPER, are designed to replace tokens in given texts by sampling from the LLM. STA-M effectively increases the proportion of green-list tokens by raising their probability in high-entropy scenarios without compromising too much text quality, making it difficult for LLM-based attacks to replace a substantial number of tokens in STA-M-generated text and remove the watermark. In conclusion, STA-M can generate text with high watermark strength against various attacks.
@@ 6 Conclusions and Future Work
In this work, we propose a new unbiased watermarking method named STA-1. We clarify the text quality (regarding the risk of unsatisfactory outputs under the same expectation) and watermark strength tradeoff of unbiased watermarks in low-entropy scenarios. We also extend STA-1 to STA-M which can enhance watermark strength with small text quality shifts. Experimental results on low-entropy datasets prove that STA-1 is comparable to other unbiased watermarks and has a low risk. Moreover, results from the high-entropy dataset demonstrate the efficiency of STA-1 and the robustness of STA-M. Future work of our study can be conducted in several ways. First, watermarking low-entropy tasks is still challenging and future work can devise better watermarking methods. Second, future work could incorporate more datasets and generative LLMs for evaluation of our method. Third, it is also possible to consider context code history to extend the unbiased results from the token level to the sequence level.
@@ 7 Related Work
With the development of LLMs, the idea of watermarking LLMs has been proposed (Aaronson, 2022; Kirchenbauer et al., 2023a) and widely explored. Existing white-box watermarking techniques can be categorized into watermarking during logits and probabilities generation (Kirchenbauer et al., 2023b; Lee et al., 2023; Hu et al., 2024; Wang et al., 2023; Fernandez et al., 2023; Zhao et al., 2023; Yoo et al., 2023; Ren et al., 2023; Takezawa et al., 2023), and watermarking by controlling sampling strategies (Christ et al., 2023; Kuditipudi et al., 2023; Hou et al., 2023; Fairoze et al., 2023). We refer readers to Appendix G for a detailed discussion on related work.
Appendix A Research Gap Summary
Table 5: Research Gap Summary. Black-box refers to only black-box LLM needed during detection; Efficiency refers to efficient detection; Robustness refers to the robustness against simple watermarking attacks. Guarantee refers to the statistical guarantee of type II error; Unbiased refers to the same expectation requirement in low-entropy scenarios.
Literature Black-box Efficiency Robustness Guarantee Unbiased
Kirchenbauer et al. (2023a)
Hu et al. (2024)
Wu et al. (2024)
Christ et al. (2023)
Kuditipudi et al. (2023)
Fairoze et al. (2023)
STA-1
Appendix B Details of Previous Methods
@@
KGW (Kirchenbauer et al., 2023a) first randomly splits the vocabulary set $\mathcal{V}$ into two non-overlapping lists based on a uniformly distributed key $k$: a ‘green’ list and a ‘red’ list. This method has two versions: the ‘hard’ version completely ignores the red list tokens and only samples tokens from the green list; The ‘soft’ version adds a predefined constant $\delta$ to logits of green list tokens while keeping logits of red list tokens fixed. The soft KGW reweights distribution as
@@
Wu et al. (2024) propose an unbiased reweighting method, named Dipmark. Dipmark arranges all probability masses over the vocabulary set from the original LLM output consecutively within the interval $[0,1]$ and then randomly permutes their orders based on a key $k$. A hyperparameter $\alpha\in[0,0.5]$ partitions the probability interval $[0,1]$ into three segments: $[0,\alpha]$, $(\alpha,1-\alpha]$, and $(1-\alpha,1]$. Probability masses in the first segment are set to 0, those in the second remain constant, and those in the third are doubled. Denote the token order after permutation as $\widetilde{\mathcal{V}}$, the adjusted probability for the $j$-th token within $\widetilde{\mathcal{V}}$ is $P_{M,w}(x^{t}=j|x^{-N_{p}:(t-1)};k)=F(j|\widetilde{\mathcal{V}})-F(j-1|\widetilde{\mathcal{V}})$, with $F(j|\widetilde{\mathcal{V}})$ being defined as
@@
Another unbiased reweighting method, RDW (robust distortion-free watermark), is developed by Kuditipudi et al. (2023). We focus on the RDW method with an inverse transform sampling scheme. In RDW, the uniformly random key $k=(\Pi,u)$, where $\Pi$ represents a random shuffle of all probability masses $P_{M}(x^{t}|x^{-N_{p}:(t-1)})$ over the vocabulary set within the interval $[0,1]$, and $u$ is a random value following the distribution $\text{U}(0,1)$. RDW first permutes the order of all $P_{M}(x^{t}|x^{-N_{p}:(t-1)})$ within the interval $[0,1]$ according to $\Pi$, then it utilizes $u$ as the cumulative distribution function value of $P_{M}(x^{t}|x^{-N_{p}:(t-1)})$ with respect to the permutation. Let $\Pi(j)$ denote the $j$-th token in the ordered vocabulary set under the permutation $\Pi$. Following the inverse transform sampling scheme, the value $u$ is inverse transformed to generate a token through
$x^{s}=\Pi(\min\{j:\;\sum_{i=1}^{j}P_{M}(x^{t}=\Pi(i)|x^{-N_{p}:(t-1)})\geq u\}),$
where $x^{s}$ is the sampled token. Therefore, we have $P_{M,w}(x^{t}=x^{s}|x^{-N_{p}:(t-1)};k)=1$, and the probabilities of all other tokens are reweighted to 00 accordingly.
Appendix C Proofs
C.1 Proof of Theorem 1
@@
C.2 Proof of Theorem 2
In this proof, we employ the notations introduced in the proof of Theorem 1 in Section C.1, and we leverage the results derived from that theorem’s proof.
@@ C.2 Proof of Theorem 2
C.3 Proof of Corollary 1
For the $z$-test in detecting STA-1, its type II error is defined as $P(z\leq\tilde{z}|H_{a})$. Following the definition, we have that
$P(z\leq\tilde{z}|H_{a})$ $=P\left(\frac{|S|_{G}-\gamma T}{\sqrt{\gamma(1-\gamma)T}}\leq\tilde{z}\bigg{|}H_{a}\right)$
$=P(|S|_{G}-\mathbb{E}(|S|_{G})\leq\gamma T+\tilde{z}\sqrt{\gamma(1-\gamma)T}-\mathbb{E}(|S|_{G})|H_{a})$
$\leq P(|S|_{G}-\mathbb{E}(|S|_{G})\leq\gamma T+\tilde{z}\sqrt{\gamma(1-\gamma)T}-\underline{\mathbb{E}}|H_{a})$
[…]
$\leq\frac{\overline{\mathbb{V}}}{\overline{\mathbb{V}}+(\underline{\mathbb{E}}-\gamma T-\tilde{z}\sqrt{\gamma(1-\gamma)T})^{2}},$
where Cantelli’s inequality holds because
$\underline{\mathbb{E}}-(\gamma T+\tilde{z}\sqrt{\gamma(1-\gamma)T})=\gamma(1-\gamma)TGini^{*}-\tilde{z}\sqrt{\gamma(1-\gamma)T}>0$
according to the condition assumed in the corollary. This completes the proof.
C.4 Proof of Theorem 3
In this proof, we continue utilizing the notations introduced in the proof of Theorem 1 in Section C.1.
@@ C.4 Proof of Theorem 3
Finally, we determine the variance for the RDW method with an inverse transform sampling scheme. In RDW, the uniformly distributed key $k=(\Pi,u)$, where $\Pi$ is a uniformly random permutation of the $N$ tokens and $u\sim\text{U}(0,1)$. Similar to the previous analysis of Dipmark, the relative orders among the remaining $(N-1)$ tokens except $x_{max}$ are irrelevant to the permutation. Therefore, we only need to consider the $N$ unique permutations, each with a probability of $1/N$, as discussed above. Conditional on any permutation $\Pi$, under the inverse transform sampling scheme, there is a probability of $p_{max}$ that $x_{max}$ will be sampled out. Therefore, the altered value of $p_{max}$ given $\Pi$ is
$p_{max}^{w,k}|\Pi=\begin{cases}1&\text{with probability }p_{max}\\ 0&\text{with probability }1-p_{max}\end{cases}.$
Then, we have that
$\mathbb{V}_{u}^{\text{RDW}}\left[p_{max}^{w,k}|\Pi\right]=p_{max}(1-p_{max}).$
Because these results hold for any permutation $\Pi$, by the law of total variance, we can derive that
$\mathbb{V}_{k\sim P_{K}(k)}^{\text{RDW}}\left[p_{max}^{w,k}\right]$ $=\mathbb{E}_{\Pi}\left(\mathbb{V}_{u}^{\text{RDW}}\left[p_{max}^{w,k}|\Pi\right]\right)+\mathbb{V}_{\Pi}\left(\mathbb{E}_{u}^{\text{RDW}}\left[p_{max}^{w,k}|\Pi\right]\right)$
$=p_{max}(1-p_{max})+0$
$=p_{max}(1-p_{max}),$
which is the variance for the RDW method with an inverse transform sampling scheme.
@@
$\mathbb{V}_{k\sim P_{K}(k)}^{\text{STA-1}}\left[p_{max}^{w,k}\right]<(1-p_{max})^{2}\frac{(N+1)}{3(N-1)}=\mathbb{V}_{k\sim P_{K}(k)}^{\text{Dipmark}}\left[p_{max}^{w,k}\right].$
For the comparison between $\mathbb{V}_{k\sim P_{K}(k)}^{\text{Dipmark}}\left[p_{max}^{w,k}\right]$ and $\mathbb{V}_{k\sim P_{K}(k)}^{\text{RDW}}\left[p_{max}^{w,k}\right]$, we have that
$\mathbb{V}_{k\sim P_{K}(k)}^{\text{Dipmark}}\left[p_{max}^{w,k}\right]$ $=(1-p_{max})^{2}\frac{(N+1)}{3(N-1)}$
$<(1-p_{max})^{2}$
$\leq p_{max}(1-p_{max})=\mathbb{V}_{k\sim P_{K}(k)}^{\text{RDW}}\left[p_{max}^{w,k}\right],$
where the first inequality holds because $(N+1)<3(N-1)$ for $N>2$, and the second inequality is valid under the assumption that $11-\alpha\leq p_{max}<$ 1 and $\alpha\in[0,0.5]$.
Putting all the results together, we get
$\mathbb{V}_{k\sim P_{K}(k)}^{\text{STA-1}}\left[p_{max}^{w,k}\right]<\mathbb{V}_{k\sim P_{K}(k)}^{\text{Dipmark}}\left[p_{max}^{w,k}\right]=\mathbb{V}_{k\sim P_{K}(k)}^{\gamma\text{-reweight}}\left[p_{max}^{w,k}\right]<\mathbb{V}_{k\sim P_{K}(k)}^{\text{RDW}}\left[p_{max}^{w,k}\right],$
@@
Appendix D Example of Risk-averse
St. Petersburg paradox (Wikipedia, 2024). Assume that one must choose either one lottery from the following two lotteries. (1) Lottery 1 ($L1$) has a 0.8 probability of earning nothing and the other 0.2 probability of losing 1,000 dollars. (2) Lottery 2 ($L2$) has a 0.5 probability of losing 100 dollars and the other 0.5 probability of losing 300 dollars.
@@ Appendix D Example of Risk-averse
Computationally, assume the person has 1,001 dollars in total and the utility function is $\ln(Y)$ (Debreu et al., 1954), where $Y$ is the wealth. The utility function measures happiness. It is a concave function (such as $\ln(Y)$) because people are happier if they are wealthier ($\ln^{\prime}(Y)>0$) but the increment of happiness decreases as the wealth increases ($\ln^{\prime\prime}(Y)<0$).
@@ Appendix E STA-M Details
Appendix F Experiment
F.1 Experimental Setup
Datasets and metrics. We employed two public datasets which are C4 subset (Raffel et al., 2020; Kirchenbauer et al., 2023a) for news-like text generation and HumanEval (Chen et al., 2021) for code generation. Specifically, C4 represents the high-entropy generation task and HumanEval represents the low-entropy generation task.
C4: We extracted random text segments from the news-like subset of the C4 dataset (Raffel et al., 2020) following Kirchenbauer et al. (2023a). For each segment, we removed a fixed number of tokens from the end and the removed tokens served as a ‘baseline’ completion. The remaining tokens were used as the prompt.